# Think Outside the Policy: In-Context Steered Policy Optimization

### v1

@@ Abstract
Existing Reinforcement Learning from Verifiable Rewards (RLVR) methods, such as Group Relative Policy Optimization (GRPO), have achieved remarkable progress in improving the reasoning capabilities of Large Reasoning Models (LRMs). However, they exhibit limited exploration due to reliance on on-policy rollouts where confined to the current policy’s distribution, resulting in narrow trajectory diversity. Recent approaches attempt to expand policy coverage by incorporating trajectories generated from stronger expert models, yet this reliance increases computational cost and such advaned models are often inaccessible. To address these issues, we propose In-Context Steered Policy Optimization (ICPO), a unified framework that leverages the inherent in-context learning capability of LRMs to provide expert guidance using existing datasets. ICPO introduces Mixed-Policy GRPO with Implicit Expert Forcing, which expands exploration beyond the current policy distribution without requiring advanced LRM trajectories. To further stabilize optimization, ICPO integrates Expert Region Reject Sampling to filter unreliable off-policy trajectories and Annealed Expert-Bonus Reward Shaping to balance early expert guidance with later autonomous improvement. Results demonstrate that ICPO consistently enhances reinforcement learning performance and training stability on mathematical reasoning benchmarks, revealing a scalable and effective RLVR paradigm for LRMs.
@@
Hsiu-Yuan Huang1,2††thanks: Equal contribution. Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3††thanks: Corresponding author., Saiyong Yang3, Yunfang Wu1,222footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University 3LLM Department, Tencent ✉ huang.hsiuyuan@stu.pku.edu.cn, wuyf@pku.edu.cn
@@ 1 Introduction
Large Reasoning Models (LRMs) excel at solving complex mathematical problems, and Reinforcement Learning from Verifiable Rewards (RLVR) provides a scalable way to refine their reasoning through symbolic correctness signals. Yet, limited exploration under standard Group Relative Policy Optimization (GRPO) Shao et al. (2024) often hinders robust reasoning improvement.
@@
To address these limitations, recent work has explored combining Supervised Fine-Tuning (SFT) with Reinforcement Learning (RL) to strengthen the exploration of LRMs. One research direction interleaves SFT and RL updates Ma et al. (2025), enabling SFT to improve high-difficulty problem-solving while RL refines mid- to low-difficulty behaviors. However, repeated switching between paradigms introduces instability and inefficient convergence. Another line of work seeks to unify SFT and RL within a single training process through different strategies: incorporating SFT data as off-policy rollouts during RL Yan et al. (2025) to expand the exploration space; jointly optimizing SFT and RL objectives Fu et al. (2025) for tighter integration; and leveraging hints to bootstrap rollouts Liu et al. (2025); Zhang et al. (2025); Fu et al. (2025) to improve performance on harder prompts.
@@
Together, these approaches reflect a growing consensus that combing SFT and RL can effectively expand the exploration of policy search into more promising reasoning spaces. However, as shown in Figure 1, three key challenges remain: (1) Exploration remains confined to the current policy distribution, as GRPO-based methods rely on on-policy sampling, resulting in limited trajectory diversity and might convergence to local optima, as detailed in Section 3. (2) Expanding the exploration space with trajectories from stronger LRMs incurs high computational cost and limited accessibility, since generating additional reasoning traces from advanced models is expensive and such models are not always available. (3) External trajectories are often noisy and unstable for training, incorporating them indiscriminately as off-policy rollouts can mislead policy updates and harm convergence stability.
@@
To address these limitations, we propose In-Context Steered Policy Optimization (ICPO), a unified RL framework that exploits the LRM’s inherent In-Context Learning (ICL) capability to provide expert guidance instead of relying on external advanced LRMs. Specifically, (1) ICPO introduces Mixed-Policy GRPO with Implicit Expert Forcing (IEF), where expert-conditioned rollouts are generated through few-shot ICL guidance, enabling exploration beyond the current policy distribution and steering the model toward expert-aligned regions of the solution space. (2) To ensure reliable guidance, ICPO employs Expert Region Reject Sampling (ERRS), which filters out noisy or low-quality off-policy trajectories using verifiable reward signals, retaining only those that truly exhibit expert-aligned reasoning behavior and preventing misleading gradients from contaminating policy updates. (3) ICPO further incorporates an annealed expert bonus into the Reward Shaping (RS) design, enforcing strong expert-guided shaping in the early stage and gradually relaxing it to facilitate autonomous optimization as exploration capabilities grow.
@@
Our experiments show that few-shot ICL rollouts exhibit superior quality and diversity, making them effective sources of expert guidance. Extensive experiments on multiple mathematical reasoning benchmarks further demonstrate that ICPO achieves substantial performance gains over vanilla and mixed-policy GRPO methods, with maximum average improvements of up to +4.17 and +2.79 points. On out-of-distribution benchmarks, ICPO also outperforms vanilla GRPO by up to +2.37 points.
@@
We empirically show that few-shot ICL rollouts provide diverse and high-quality expert signals for mathematical reasoning task.
We propose In-Context Steered Policy Optimization, which include IEF that leverages the model’s inherent ICL ability without relying on stronger external LRMs, together with ERRS and RS for stable and efficient optimization.
ICPO delivers consistent improvements on mathematical reasoning benchmarks across model scales, demonstrating strong potential for LRM post-training.
@@ 2 Related Work
SFT and RL for LLM Reasoning.
GRPO-based RL improves reasoning by sampling and reinforcing high-quality trajectories, yet its exploration remains confined to the current policy distribution, leaving some prompts remain unsolved within feasible rollout budgets. Mainstream approaches introduce SFT to provide complementary trajectories, thereby expanding the reasoning space and enabling RL to further refine its behavior.
Several works explore hybrid SFT+RL strategies. ReLIFT Ma et al. (2025) alternates between RL and SFT by updating on failed rollouts. LUFFY Yan et al. (2025) incorporates SFT trajectories as off-policy samples using importance sampling. SRFT Fu et al. (2025) jointly optimizes SFT and RL objectives with an entropy-based weight on the SFT loss. Other approaches guide rollouts by concatenating partial SFT solutions as hints Liu et al. (2025); Zhang et al. (2025); Huang et al. (2025). However, these methods rely on advanced LRMs to supply SFT traces, which may incur additional computation overhead. In contrast, we leverage the inherent ICL ability of LRMs to obtain diverse rollouts from existing datasets, requiring neither external LRMs’ trajectories nor explicitly engineered hints.
@@ 3 Preliminary
Few-Shot ICL as Expert-Conditioned Inference.
Few-shot ICL provides a gradient-free way to inject expert priors through the input context. As illustrated in Figure 2 and Figure 3, we evaluate this effect from three complementary perspectives. (1) Accuracy: the 1-shot ICL setting consistently outperforms the 0-shot baseline, indicating that conditioning on demonstrations improves reasoning correctness. (2) Diversity: compared with temperature-based sampling, introducing 1-shot demonstrations expands the sampling space, yielding larger inter-trajectory similarity distances and enhanced exploratory diversity. (3) Distribution quality: under ICL-conditioned rollouts, the output distribution becomes more favorable—a higher proportion of previously incorrect generations are “flipped” to correct solutions compared with temperature perturbations, indicating that in-context steering provides a stronger and more targeted exploration signal. Taken together, these results support our view that few-shot ICL constitutes an effective expert-conditioned inference process.
From Few-Shot ICL to Implicit Expert Forcing.
@@
indicating that ICL implicitly introduces an expert-induced prior that steers the rollout distribution toward expert-like regions—without any explicit optimization on . While few-shot ICL itself is an inference-time mechanism that does not update model parameters, we incorporate its induced trajectories into RL training to form IEF.
@@
GRPO is an efficient On-Policy optimization algorithm tailored for RL in LLMs, where the advantages for each token are computed in a group-relative manner without requiring an additional critic model to estimate token values. Given a set of rollouts sampled from the old policy , the normalized advantage is computed by:
@@ 4.1 Mixed-Policy GRPO with Implicit Expert Forcing
The objective of mixed-policy GRPO with IEF can be written as follows:
where is the expert-conditioned importance weight. The mixed objective balances exploitation within the current policy support and exploration toward expert-aligned regions that are unreachable by standard on-policy rollouts.
Unlike prior work Yan et al. (2025), which adopts a model-based off-policy scheme by relying on an additional advanced LRM to provide expert trajectories, our mixed-policy GRPO with IEF operates as an input-conditioned off-policy method. Here, in-context demonstrations steer the same policy away from its natural output distribution, producing rollouts that differ from standard on-policy samples , thus effectively forming an off-policy distribution induced by IEF.
@@ 4.2 Expert Region Reject Sampling
where normalizes over all valid tokens. The shaping function follows prior work Yan et al. (2025) and is defined as , where by default. This shaping biases learning toward expert-induced improvements while encouraging exploration.
@@ 4.3 Reward Shaping with Annealed Expert Bonus
[…]
@@
To encourage early imitation of expert-conditioned behavior while avoiding long-term over-reliance, we add a step-annealed bonus only to trajectories that have correct answer and within the expert region :
@@ 5 Experimental Setup
We follow Yan et al. (2025) and adopt the OpenR1-Math-220k dataset as our main training corpus. This dataset comprises mathematical problems sourced from NuminaMath-1.5 Li et al. (2024) and reasoning traces generated by the advanced LRM DeepSeek-R1 DeepSeek-AI (2025). Specifically, we use the filtered subset111https://huggingface.co/datasets/Elliott/Openr1-Math-46k, which excludes generations exceeding 8192 tokens as well as those identified as incorrect by Math-Verify222https://github.com/huggingface/Math-Verify. The resulting dataset contains approximately 45k verified reasoning prompts.
For the ablation study, we additionally train on a simpler dataset, Skywork-OR1-RL-Data333https://huggingface.co/datasets/Skywork/OR1-RL-Data He et al. (2025), to verify the generalization ability of our IEF under different reasoning conditions. This dataset is annotated with difficulty levels predicted by DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025), and we use the subset with difficulty level = 1, which corresponds to the easiest reasoning problems, as our simplified training corpus for controlled comparison.
@@
We use Qwen3-1.7B and Qwen3-8B Yang et al. (2025) as the base model and employ GRPO Shao et al. (2024) as our RL algorithm. The KL regularization term is retained, while the entropy coefficient is set to 0.0 to stabilize training. We use a rollout batch size of 128 and an update batch size of 64. We generate 8 rollout trajectories per prompt with a maximum sequence length of 8192 tokens. Specifically, for the on-policy baseline, we use 8 on-policy rollouts. For our mixed-policy GRPO, we follow previous work Yan et al. (2025) and use 1 off-policy and 7 on-policy rollouts to ensure comparability. For implicit expert forcing, we randomly sample demonstrations from the MATH Hendrycks et al. (2021) training set, which contains 7.5k mathematical problems paired with high-quality solutions. Rollout generation uses a temperature of 1.0, and rewards are computed using Math-Verify. All models are trained for optimization steps, and we report results using the final checkpoint. Details of our evaluation setting are provided in the Appendix B.
@@
For RLVR-based methods, we compare against the vanilla GRPO baseline, which does not incorporate any external expert knowledge. We also include LUFFY Yan et al. (2025), which leverages trajectories generated by advanced LRMs as off-policy rollouts.
6 Results and Analysis
@@ 6.1 Main Results
| Expert Domain | In-Distribution Benchmarks | Overall | |||||||
@@
| Model | Method | MATH-500 | AIME24 | AIME25 | AMC23 | Minerva | Olympiad | Avg. (Impr.) | |
| Qwen3-1.7B | GRPO | 83.60 | 28.44 | 22.50 | 66.72 | 40.81 | 48.15 | 48.37 | |
| ICPO | 86.80 | 31.25 | 26.25 | 70.39 | 44.12 | 56.44 | 52.54 | (+4.17) | |
| ICPO | 87.20 | 28.96 | 26.56 | 70.00 | 42.65 | 52.74 | 51.35 | (+2.98) | |
| Qwen3-8B | GRPO | 91.00 | 54.79 | 38.54 | 83.75 | 50.74 | 62.37 | 63.53 | |
| ICPO | 92.00 | 55.21 | 43.65 | 86.95 | 51.10 | 65.19 | 65.68 | (+2.15) | |
| ICPO | 92.00 | 56.15 | 40.94 | 92.00 | 51.47 | 64.30 | 65.04 | (+1.51) | |
| Method | ARC | GPQA | MMLU | Avg. (Impr.) |
| Qwen3-1.7B | ||||
| GRPO | 88.31 | 34.34 | 54.43 | 59.03 |
| ICPO | 88.14 | 27.78 | 55.45 | 57.12 (-1.91) |
| ICPO | 87.71 | 36.36 | 54.95 | 59.67 (+0.64) |
| Qwen3-8B | ||||
| GRPO | 95.82 | 51.01 | 71.98 | 72.94 |
| ICPO | 95.48 | 55.05 | 72.30 | 74.28 (+1.34) |
| ICPO | 95.56 | 55.05 | 75.31 | 75.31 (+2.37) |
@@
The main experiments include two variants of our proposed ICPO framework: ICPO, which operates without reward shaping (RS), and ICPO, which incorporates RS to further enhance expert-domain alignment. To better understand their optimization behavior against GRPO, we further visualize the reward dynamics over training steps across different datasets, as shown in Figure 5, where both ICPO variants consistently achieve higher rewards throughout training.
Overall Improvements.
@@
Table 1 reports the in-distribution reasoning performance, where MATH-500 serves as the expert domain. Across both model scales, ICPO consistently outperforms the vanilla GRPO baseline. For the smaller Qwen3-1.7B model, ICPO achieves an average improvement of +2.98 points over GRPO, while ICPO further stabilizes optimization with a +4.17 point overall gain. A similar trend is observed for the larger Qwen3-8B model, where ICPO and ICPO yield +2.15 and +1.51 average improvements, respectively. These consistent gains across scales demonstrate that our ICPO effectively steers the policy toward more expert-aligned regions in the parameter space.
Effect of Reward Shaping.
The variant ICPO introduces RS to explicitly amplify the advantage of trajectories falling within the expert domain . This design encourages the model to allocate higher probability mass to expert-like reasoning trajectories. Consequently, ICPO achieves stronger performance on the expert domain dataset (MATH-500), yielding +3.6 point improvements for Qwen3-1.7B. The improvement is less pronounced in non-expert domains, suggesting that RS primarily enhances expert-domain alignment rather than generic exploration. We note that the performance of ICPO could be further enhanced with higher-quality expert data, which would allow more precise shaping of expert-domain rewards and stronger alignment effects.
@@
To assess whether ICPO preserves the model’s general reasoning capabilities, we further evaluate it on several OOD benchmarks, as shown in Table 2. The results indicate that ICPO maintains competitive or superior generalization performance compared to GRPO across both model scales, especially on larger model. For Qwen3 1.7B and 8B, ICPO achieves the average improvement of +0.64 and +2.37 over GRPO, suggesting that moderate RS enhances domain-specific reasoning while preserving broad OOD generalization. We also note that smaller models exhibit less stable OOD generalization than larger ones, likely due to their limited reasoning capacity.
Comparison of Expert Guidance Sources.
| Method | A. | E. | R. | MATH | AIME24/25 | AMC | Mnrv. | Avg. |
| Direct Comparison of Expert Guidance Sources | ||||||||
| LUFFY | ✓ | ✗ | ✗ | 91.00 | 53.12 / 36.98 | 85.31 | 52.21 | 63.72 |
| ICPO∗ | ✗ | ✓ | ✗ | 89.60 | 55.21 / 41.67 | 85.16 | 53.31 | 64.99 |
| In-Context Steered Policy Optimization | ||||||||
| ICPO | ✗ | ✓ | ✓ | 92.00 | 55.21 / 43.65 | 86.95 | 51.10 | 65.78 |
| ICPO | ✗ | ✓ | ✓ | 92.00 | 56.15 / 40.94 | 92.00 | 51.47 | 66.51 |
We compare ICPO with LUFFY Yan et al. (2025), which incorporates trajectories generated by advanced LRMs into off-policy GRPO, as shown in Table 3. ICPO∗, which removes ERRS and is thus directly comparable to LUFFY in its source of expert guidance, already surpasses LUFFY by an average of +1.27 points. This demonstrates that IEF can steer the model toward a better policy distribution by leveraging existing datasets as contextual guidance, eliminating the need for costly external LRM computation. Moreover, incorporating ERRS and RS further improves performance by +2.06 and +2.79 points over LUFFY, confirming their complementary benefits.

### v2

@@ Abstract
Existing Reinforcement Learning from Verifiable Rewards (RLVR) methods, such as Group Relative Policy Optimization (GRPO), have achieved remarkable progress in improving the reasoning capabilities of Large Reasoning Models (LRMs). However, they exhibit limited exploration due to reliance on on-policy rollouts which are confined to the current policy’s distribution, resulting in narrow trajectory diversity. Recent approaches attempt to expand policy coverage by incorporating trajectories generated from stronger expert models, yet this reliance increases computational cost and such advanced models are often inaccessible. To address these issues, we propose In-Context Steered Policy Optimization (ICPO), a unified framework that leverages the inherent in-context learning capability of LRMs to provide expert guidance using existing datasets. ICPO introduces mixed-policy GRPO with implicit expert forcing, which expands exploration beyond the current policy distribution without requiring advanced LRM trajectories. To further stabilize optimization, ICPO integrates expert region reject sampling to filter unreliable off-policy trajectories and annealed expert-bonus reward shaping to balance early expert guidance with later autonomous improvement. Results demonstrate that ICPO consistently enhances RLVR performance and training stability on mathematical reasoning benchmarks, revealing a scalable and effective RLVR paradigm for LRMs. Our code is available at https://anonymous.4open.science/r/ICPO.
@@
Hsiu-Yuan Huang1,2††thanks: Equal contribution. Work done during internship at Tencent., Chenming Tang1,211footnotemark: 1, Weijie Liu3††thanks: Corresponding author., Clive Bai3, Saiyong Yang3, Yunfang Wu1,222footnotemark: 2 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Computer Science, Peking University 3LLM Department, Tencent ✉ huang.hsiuyuan@stu.pku.edu.cn, wuyf@pku.edu.cn
@@ 1 Introduction
Large Reasoning Models (LRMs) excel at solving complex mathematical problems, and Reinforcement Learning from Verifiable Rewards (RLVR) provides a scalable way to refine their reasoning through verifiable rewards DeepSeek-AI (2025). Yet, limited exploration under standard Group Relative Policy Optimization (GRPO) Shao et al. (2024) often hinders robust reasoning improvement Yue et al. (2025); Zhang et al. (2025a).
@@
Recent work has explored combining Supervised Fine-Tuning (SFT) with Reinforcement Learning (RL) to strengthen the exploration of LRMs. One research direction interleaves SFT and RL updates Ma et al. (2025), enabling SFT to improve high-difficulty problem-solving while RL refines mid- and low-difficulty behaviors. However, repeated switching between paradigms introduces instability and inefficient convergence. Another line of work seeks to unify SFT and RL within a single training process through different strategies: incorporating SFT data as off-policy rollouts during RL to expand the exploration space Yan et al. (2025); jointly optimizing SFT and RL objectives for tighter integration Fu et al. (2025); and leveraging hints to bootstrap rollouts to improve performance on harder prompts Liu et al. (2025a); Zhang et al. (2025b); Fu et al. (2025).
@@
Together, these approaches reflect a growing consensus that combining SFT and RL can effectively expand the exploration of policy search into more promising reasoning spaces. However, as shown in Figure 1, three key challenges remain: (1) Online exploration remains confined to the current policy distribution, as GRPO-based methods rely on on-policy sampling, resulting in limited trajectory diversity and might converge to local optima. (2) Expanding the exploration space with trajectories from stronger LRMs incurs high computational cost and limited accessibility, since generating additional reasoning traces from advanced models is expensive and such models are not always available. (3) External trajectories are often noisy and unstable for training, incorporating them indiscriminately as off-policy rollouts can mislead policy updates and harm convergence stability.
@@
To address these limitations, we propose In-Context Steered Policy Optimization (ICPO), a novel RLVR framework that exploits the LRM’s inherent In-Context Learning (ICL) capability to provide expert guidance instead of relying on external advanced LRMs. Specifically, (1) ICPO introduces mixed-policy GRPO with Implicit Expert Forcing (IEF), where expert-conditioned rollouts are generated through ICL guidance, enabling exploration beyond the current policy distribution and steering the model toward expert-aligned regions of the solution space. (2) To ensure reliable guidance, ICPO employs Expert Region Reject Sampling (ERRS), which filters out noisy or low-quality off-policy trajectories using verifiable reward signals, preventing misleading gradients from contaminating policy updates. (3) ICPO further explores an annealed expert bonus into the Reward Shaping (RS) design, enforcing strong expert-guided shaping in the early stage and progressively relaxing it to facilitate autonomous optimization as LRM capabilities grow.
@@
Our experiments show that ICL-based rollouts exhibit superior quality and diversity, making them effective sources of expert guidance (§ 3). Extensive evaluations across multiple mathematical reasoning benchmarks further demonstrate that ICPO achieves substantial performance gains over vanilla GRPO and mixed-policy GRPO Yan et al. (2025), with maximum average improvements of up to +4.1 and +4.0 points, respectively (§ 6). In addition, we provide in-depth instance- and token-level analyses that offer strong evidence of distributional shifts induced by ICPO (§ 7).
@@
We empirically show that ICL rollouts provide diverse and high-quality expert signals for the mathematical reasoning task.
We introduce ICPO, which includes mixed-policy GRPO with Implicit Expert Forcing, leveraging the model’s inherent ICL ability without relying on stronger LRMs, together with Expert Region Reject Sampling and Reward Shaping for stable and efficient learning.
We empirically validate that ICPO delivers consistent improvements on mathematical reasoning benchmarks across model scales and effectively encourages exploration, demonstrating strong potential for LRM post-training.
@@ 2 Related Work
Exploration from Within the Policy.
Recent methods enhance exploration by increasing diversity within the policy itself, such as generating additional rollouts Wang et al. (2025); Hu et al. (2025a); Tang et al. (2025), leveraging replay buffers to revisit informative prompts Dou et al. (2025), and sampling under higher temperatures to boost stochasticity Chen et al. (2025); Zhang et al. (2024). These approaches broaden trajectory coverage under the model’s own distribution. However, their exploration remains bounded by the intrinsic capacity of the current policy.
Exploration from Outside the Policy.
Other studies leverage external supervision to guide exploration beyond the model’s default policy distribution. One line of work explores hybrid SFT–RL strategies to expand the reasoning space of LRMs. ReLIFT Ma et al. (2025) alternates between RL and SFT by updating on failed rollouts. LUFFY Yan et al. (2025) incorporates SFT trajectories as off-policy samples using importance sampling. SRFT Fu et al. (2025) jointly optimizes SFT and RL objectives with an entropy-based weight on the SFT loss. Another line of work guides rollouts by concatenating partial SFT solutions as hints Liu et al. (2025a); Zhang et al. (2025b); Huang et al. (2025). However, these methods rely on advanced LRMs to supply SFT traces, which may incur additional computation overhead.
In contrast, we exploit LRMs’ inherent ICL ability to steer diverse rollouts with existing datasets as demonstrations, requiring neither external LRMs’ trajectories nor explicitly engineered hints.
@@ 3 Preliminary
ICL as Expert-Conditioned Inference.
ICL provides a gradient-free way to inject expert priors through the input context, which we evaluate from three complementary perspectives. (1) Accuracy: the 1-shot ICL setting consistently outperforms the 0-shot baseline (Figure 2), indicating that conditioning on demonstrations improves reasoning correctness. (2) Diversity: compared with temperature-based sampling, introducing 1-shot demonstrations expands the sampling space (Figure 3), yielding larger inter-trajectory edit distances and enhanced exploratory diversity. (3) Distribution quality: under ICL-conditioned rollouts, the output distribution becomes more favorable—a higher proportion of previously incorrect generations are “flipped” to correct solutions compared with temperature perturbations (Figure 3), indicating that in-context steering provides a stronger and more targeted exploration signal. Taken together, these results support our view that ICL constitutes an effective expert-conditioned inference process.
From ICL to Implicit Expert Forcing.
@@
indicating that ICL implicitly introduces an expert-induced prior that steers the rollout distribution toward expert-like regions—without any explicit optimization on . While ICL itself is an inference-time mechanism that does not update model parameters, we incorporate its induced trajectories into GRPO training to form IEF.
@@
GRPO is an efficient On-Policy optimization algorithm tailored for RL in LLMs, where the advantage for each token is computed in a group-relative manner without requiring an additional critic model to estimate token values. Given a set of rollouts sampled from the old policy , the normalized advantage is computed by:
@@ 4.1 Mixed-Policy GRPO with Implicit Expert Forcing
The objective of mixed-policy GRPO with IEF balances exploitation of the current policy with exploration toward expert-aligned regions, and can be written as follows:
where is the expert-conditioned importance weight.
Unlike prior work Yan et al. (2025), which adopts a model-based off-policy scheme by relying on an additional advanced LRM to provide expert trajectories for the same training prompts, our mixed-policy GRPO with IEF operates as an input-conditioned off-policy method. Specifically, while all rollouts are sampled from the same policy , in-context demonstrations alter the input conditioning and steer the policy away from its default output distribution. This conditioning discrepancy induces a behavior mismatch, under which expert-conditioned rollouts are regarded as off-policy relative to standard on-policy samples .
@@ 4.2 Expert Region Reject Sampling
where normalizes over all valid tokens. The shaping function follows prior work and is defined as , where by default Yan et al. (2025). This shaping biases learning toward expert-induced improvements while encouraging exploration.
@@ 4.3 Reward Shaping with Annealed Expert Bonus
The verifiable reward function evaluates the model output by extracting the final answer and comparing it against the predefined ground-truth answer. It assigns a binary score based on whether the extracted answer matches the correct solution under a task-specific verifier. Formally,
@@
To encourage early imitation of expert-conditioned behavior while avoiding long-term over-reliance, we design a variant of ICPO (namely ICPO) and add a step-annealed bonus only to trajectories that have the correct answer and within the expert region :
@@ 5 Experimental Setup
We follow Yan et al. (2025) and adopt the OpenR1-Math-220k dataset as our main training corpus. Specifically, we use the filtered subset111https://huggingface.co/datasets/Elliott/Openr1-Math-46k, which excludes generations exceeding 8192 tokens as well as those identified as incorrect by Math-Verify222https://github.com/huggingface/Math-Verify. The resulting dataset contains approximately 45k verified reasoning prompts. For IEF, instead of using trajectories from advanced LRMs, we randomly sample demonstrations for each prompt from the MATH Hendrycks et al. (2021) training set, which contains 7.5k mathematical problems paired with high-quality solutions.
@@
The complete set of hyperparameters used in our training is listed in Appendix B.1. For the main experiments, we use Qwen3-1.7B and Qwen3-8B Yang et al. (2025) as the base models and employ GRPO Shao et al. (2024) as our RL algorithm. We also include Qwen2.5-Math-7B to facilitate comparison with other related methods. We additionally evaluate ICPO on Qwen3-8B-Base and LLaMA-3.1-8B to examine its cross-model generalization. We generate a total of 8 rollout trajectories per prompt. For the on-policy baseline, we use 8 on-policy rollouts. For our mixed-policy GRPO, we follow previous work Yan et al. (2025) and use 1 off-policy rollout and 7 on-policy rollouts to ensure comparability. Evaluation details are provided in Appendix B.2.
@@
For direct comparison, we evaluate our ICPO against the vanilla GRPO baseline trained on the same subset of OpenR1-Math-220k. We further include , which extends GRPO by using 16 on-policy rollouts to enhance exploration. To examine the effect of training data source, we also train a GRPO variant on expert-domain data, identical to our IEF demonstrations (MATH), denoted as . In addition, we compare with LUFFY Yan et al. (2025), which leverages trajectories generated by advanced LRMs as off-policy rollouts, along with other baselines detailed in Appendix B.3.
6 Experimental Results
@@ 6.1 Main Results
| In-Distribution Benchmarks | Out-of-Distribution Benchmarks | |||||||||||
@@
| Method | AIME 24/25 | AMC23 | MATH | Minerva | Olympiad | Avg. | (Impr.) | ARC | GPQA | MMLU | Avg. | (Impr.) |
| Qwen3-1.7B | 21.7 / 20.6 | 56.8 | 79.4 | 37.1 | 40.6 | 42.7 | – | 88.3 | 22.2 | 52.3 | 54.3 | – |
| GRPO | 28.4 / 22.5 | 66.7 | 83.6 | 40.8 | 48.2 | 48.4 | – | 88.3 | 34.3 | 54.4 | 59.0 | – |
| 28.3 / 24.7 | 69.8 | 84.4 | 43.0 | 53.8 | 50.7 | (+2.3) | 88.9 | 30.8 | 55.0 | 58.2 | (-0.8) | |
| 26.8 / 24.8 | 66.3 | 83.8 | 45.6 | 50.5 | 49.6 | (+1.2) | 88.8 | 32.8 | 55.5 | 59.0 | (0.0) | |
| ICPO (Ours) | 31.3 / 26.3 | 70.4 | 86.8 | 44.1 | 56.4 | 52.5 | (+4.1) | 88.1 | 27.8 | 55.5 | 57.1 | (-1.9) |
| ICPO (Ours) | 29.0 / 26.6 | 70.0 | 87.2 | 42.7 | 52.7 | 51.4 | (+3.0) | 87.7 | 36.4 | 55.0 | 59.7 | (+0.7) |
| Qwen3-8B | 33.7 / 21.4 | 65.4 | 86.2 | 40.4 | 43.3 | 48.4 | – | 96.2 | 37.9 | 68.0 | 67.4 | – |
| GRPO | 54.8 / 38.5 | 83.8 | 91.0 | 50.7 | 62.4 | 63.5 | – | 95.8 | 51.0 | 72.0 | 72.9 | – |
| 53.0 / 40.0 | 85.4 | 93.0 | 52.9 | 61.6 | 64.3 | (+0.8) | 95.7 | 53.0 | 72.9 | 73.9 | (+1.0) | |
| 58.3 / 40.4 | 86.2 | 91.8 | 50.7 | 60.0 | 64.6 | (+1.1) | 92.2 | 51.0 | 70.2 | 71.1 | (-1.8) | |
| ICPO (Ours) | 55.2 / 43.7 | 87.0 | 92.0 | 51.1 | 65.2 | 65.7 | (+2.2) | 95.5 | 55.1 | 72.3 | 74.3 | (+1.4) |
| ICPO (Ours) | 56.2 / 40.9 | 85.4 | 92.0 | 51.5 | 64.3 | 65.0 | (+1.5) | 95.6 | 55.1 | 75.3 | 75.3 | (+2.4) |
| Method | AIME24 | AIME25 | AMC23 | MATH | Minerva | Olympiad | Average |
| Qwen2.5-Math-7B Yang et al. (2025) | 11.5 | 4.9 | 31.3 | 43.6 | 7.4 | 15.6 | 19.0 |
| Previous RLVR Methods | |||||||
| SimpleRL-Zero* Zeng et al. (2025) | 27.0 | 6.8 | 54.9 | 76.0 | 25.0 | 34.7 | 37.4 |
| PRIME-Zero* Cui et al. (2025) | 17.0 | 12.8 | 54.0 | 81.4 | 39.0 | 40.3 | 40.8 |
| OpenReasoner-Zero* Hu et al. (2025b) | 16.5 | 15.0 | 52.1 | 82.4 | 33.1 | 47.1 | 41.0 |
| Oat-Zero* Liu et al. (2025b) | 33.4 | 11.9 | 61.2 | 78.0 | 34.6 | 43.4 | 43.8 |
| SFT and RL | |||||||
| SFT Yan et al. (2025) | 23.8 | 22.9 | 61.8 | 82.2 | 37.9 | 42.1 | 45.1 |
| RL Shao et al. (2024) | 22.9 | 13.3 | 63.0 | 81.2 | 37.1 | 43.1 | 43.4 |
| SFT+RL Shao et al. (2024) | 29.3 | 22.3 | 67.1 | 85.8 | 44.1 | 50.8 | 49.9 |
| Comparable Baseline Methods | |||||||
| ReLIFT* Ma et al. (2025) | 28.2 | 22.9 | 64.8 | 85.0 | 37.1 | 54.9 | 48.8 |
| LUFFY* Yan et al. (2025) | 29.4 | 23.1 | 65.6 | 87.6 | 37.5 | 57.2 | 50.1 |
| Prefix-RFT* Huang et al. (2025) | 31.8 | 26.4 | 68.2 | 88.4 | 40.3 | 55.7 | 51.8 |
| ICPO (Ours) | 29.3 | 28.9 | 74.9 | 88.4 | 45.6 | 52.4 | 53.2 |
| ICPO (Ours) | 32.8 | 26.7 | 75.9 | 86.6 | 44.9 | 53.6 | 53.4 |
@@
The main experiments include two variants of our proposed ICPO framework: ICPO, which operates without reward shaping (RS), and ICPO, which incorporates RS to further enhance expert-domain alignment. To better understand their optimization behavior against GRPO, we visualize the reward dynamics over training steps across different datasets, as shown in Figure 5 and 12, where both ICPO variants consistently achieve higher rewards throughout training. We also visualize the training dynamics in Appendix D.1.
In-Distribution Evaluation.
@@
Table 1 shows both the in- and out-of-distribution performance, where MATH-500 serves as the expert domain. Across both model scales, ICPO consistently outperforms the vanilla GRPO baseline, especially on in-distribution benchmarks. For the smaller Qwen3-1.7B model, ICPO achieves an average improvement of +3.0 points over GRPO, while ICPO further stabilizes optimization with a +4.1 point overall gain. A similar trend is observed for the larger Qwen3-8B model, where ICPO and ICPO yield +2.2 and +1.5 average improvements, respectively.
These consistent gains across scales indicate that ICPO effectively steers the policy toward a more optimal output distribution. By leveraging both the challenging training set and high-quality expert demonstrations through IEF, ICPO forms a more informative training signal than either source alone—a synergistic effect that yields superior performance over both and vanilla GRPO. Moreover, unlike ICPO, simply increasing the number of on-policy rollouts in offers only marginal improvement, revealing the inherent exploration limitations of vanilla GRPO and underscoring the importance of effective steering.
@@
ICPO introduces RS to explicitly amplify the advantage of trajectories falling within the expert domain . Our hypothesis is that steering optimization toward these expert-aligned trajectories induces more coherent reasoning structures and stabilizes policy updates, which in turn enhance generalization both in- and out-of-distribution. Empirically, ICPO achieves average gains of +0.7 and +2.4 over GRPO on Qwen3-1.7B and 8B, respectively, providing evidence supporting this hypothesis.
Comparison with Other Baselines.
Table 2 presents in-distribution results compared to existing baselines. ICPO achieves the strongest performance among all comparable methods, surpassing ReLIFT Ma et al. (2025), LUFFY Yan et al. (2025), and Prefix-RFT Huang et al. (2025) by +4.6, +3.3, and +1.6 points, respectively. For brevity, the results and analysis on OOD benchmarks are provided in Appendix D.3.